### v4

# Small Neutrino Masses from a Decoupled Singlet Scalar Field

Abstract
We propose a unified solution with $Z_{4}$ discrete symmetry for small neutrino masses and stability of dark matter. The Standard Model is extended with an inert doublet scalar, a dark singlet scalar, a spurion scalar and right-handed neutrinos, which all transform nontrivially under $Z_{4}$. After the $Z_{4}$ symmetry is broken to $Z_{2}$ by the VEV of the spurion, much below the mass scale of the dark singlet scalar, a small lepton number violating coupling for the inert doublet is generated at tree level, so small neutrino masses are obtained at one-loops for relatively light new fields. We discuss the important roles of the $Z_{4}$ symmetry for neutrino masses, dark matter physics and thermal leptogenesis.
I Introduction
Neutrino oscillations ParticleDataGroup:2022pth indicate neutrinos masses of $m_{\nu}\sim 0.05\,{\rm eV}$. Moreover, the bounds from neutrinoless double beta decay experiments KATRIN:2021uub and Cosmic Microwave Background anisotropies and Baryon Acoustic Oscillations Planck:2018vyg are $m_{{\bar{\nu}}_{e}}<0.8\,{\rm eV}$ and $\sum_{i}m_{\nu_{i}}<0.12\,{\rm eV}$ at 95% CL, respectively. In order to account for small neutrino masses, it is possible to extend the Standard Model (SM) by the dimension-5 operators for neutrino masses with a cutoff scale beyond the electroweak scale or light sterile neutrinos coupled to neutrinos much below the electroweak scale Cai:2017jrq .
Small neutrino masses can be obtained by the mixing between active neutrinos and heavy Majorana neutrinos Minkowski:1977sc ; Yanagida:1979as ; Gell-Mann:1979vob ; Mohapatra:1979ia . Alternatively, small neutrino masses can be generated by loop corrections, for instance, due to right-handed (RH) neutrinos and an extra doublet scalar Ma:2006km , the latter of which serves as a dark matter candidate Barbieri:2006dq ; LopezHonorez:2006gr . However, there is still a need of very small parameters in obtaining small neutrino masses in similar models Cai:2017jrq .
In this article, we propose a novel model with a $Z_{4}$ discrete symmetry for the common origin of small neutrino masses and stability of dark matter. In our model, there are one inert doublet scalar, three right-handed neutrinos, a real spurion scalar and a singlet complex scalar field, all transforming nontrivially under $Z_{4}$. The lepton number is respected in the renormalizable interactions due to $Z_{4}$. Then, only after the $Z_{4}$ symmetry is broken to $Z_{2}$ due to the VEV of the spurion field, the lepton number is broken, so neutrino masses are generated at one-loops. Moreover, the lightest neutral dark scalar in the model becomes a dark matter candidate. The spurion scalar can be regarded as a pseudo-Goldstone boson coming from the breaking of a global $U(1)$ symmetry into $Z_{4}$ at a higher scale, and $Z_{4}$ can be further down to $Z_{2}$ at a low scale.
Taking the decoupling limit where the singlet scalar field $S$ becomes heavier than the $Z_{4}$ breaking scale, we show how a small mass splitting between the neutral components of the inert doublet arises, in turn giving rise to small neutrino masses at one-loops. In the present case, we argue that large masses of extra fields or small Yukawa couplings are not necessary, unlike the case with the $Z_{2}$ symmetry only. We discuss the roles of the new singlet scalars and constraints for neutrino masses, dark matter and thermal leptogenesis.
II Discrete symmetry and dark matter scalars
We introduce a complex singlet scalar $S$ and one extra electroweak doublet scalar, $H_{2}$, other than the SM Higgs doublet $H_{1}$, and three right-handed (RH) neutrinos, $N_{R,i}$, with $i=1,2,3$, and a spurion real scalar field $\varphi$, beyond the SM. We impose a $Z_{4}$ symmetry for the extra fields such that $\phi_{i}\to e^{iQ_{X}\pi/2}\phi_{i}$ with $\phi_{i}=S,\varphi,H_{1},H_{2},N_{R}$ and $Q_{X}(\phi_{i})=(+1,+2,0,+1,+1)$. But, the SM fermions are neutral under $Z_{4}$ so the model is different from the flavor-dependent $Z_{4}$ symmetry introduced in Ref. Babu:2013pma . The $Z_{4}$ symmetry can be embedded into a global $U(1)$ Peccei-Quinn symmetry which is anomalous in the presence of the couplings to the SM quarks or extra quarks, but we don’t pursue this possibility in detail in the current work. As the subgroup of the $Z_{4}$ symmetry, we can identify $Z_{2}$ parities as $Z_{2}(\phi_{i})=(-,+,+,-,-)$. The spurion field $\varphi$ transforms by $\varphi\to-\varphi$ under the $Z_{4}$ symmetry but it is invariant under the $Z_{2}$ subgroup.
Now we consider the relevant interaction Lagrangian for the model as
${\cal L}_{\rm int}$ $=$ $-\sqrt{2}\kappa S^{\dagger}H^{\dagger}_{1}H_{2}-\sqrt{2}\lambda^{\prime}_{S\varphi}\varphi SH^{\dagger}_{1}H_{2}$ (1)
$-y_{N,ij}{\bar{l}}_{i}{\widetilde{H}}_{2}N_{R,j}-\frac{1}{2}\lambda_{N,ij}\varphi\overline{N^{c}_{R,i}}N_{R,j}+{\rm h.c.}$
with ${\widetilde{H}}_{2}=i\sigma_{2}H^{*}_{2}$. It is remarkable that $\frac{1}{2}\lambda_{5}(H^{\dagger}_{1}H_{2})^{2}$ and its hermitian conjugate, vanish due to the $Z_{4}$ symmetry, and there is no bare mass terms for right-handed neutrinos due to the lepton number conservation. There are extra $Z_{4}$ invariant interaction terms in the potential such as $\mu\varphi S^{2},\frac{1}{2}\lambda_{H_{i}\varphi}\varphi^{2}|H_{i}|^{2}$($i=1,2$), etc, as shown in the appendix, but it is sufficient to focus on eq. (1) for our discussion below.
After $\varphi$ gets a VEV by $\langle\varphi\rangle=v_{\varphi}$, $Z_{4}$ is broken to $Z_{2}$ and we obtain the $Z_{2}$ invariant effective interactions and the mass terms for right-handed neutrinos,
${\cal L}_{\rm int,eff}$ $=$ $-\sqrt{2}\kappa S^{\dagger}H^{\dagger}_{1}H_{2}-\sqrt{2}\kappa^{\prime}SH^{\dagger}_{1}H_{2}$ (2)
$-y_{N,ij}{\bar{l}}_{i}{\widetilde{H}}_{2}N_{R,j}-\frac{1}{2}M_{N,i}\overline{N^{c}_{R,i}}N_{R,i}+{\rm h.c.}$
where $\kappa^{\prime}\equiv\lambda^{\prime}_{S\varphi}v_{\varphi}$ is the $Z_{4}$ breaking effective interaction and $M_{N,i}\equiv\lambda_{N,ii}v_{\varphi}$ are the masses for RH neutrinos. Here, we assumed the basis where $\lambda_{N,ij}$ is diagonalized. After electroweak symmetry breaking with $\langle H_{1}\rangle=(0,v_{H})^{T}/\sqrt{2}\neq 0$, there appear mass mixings between dark matter scalars, namely, $S$ and the neutral components of $H_{2}$, so nonzero neutrino masses can be generated at one-loops, which will be discussed later. When the SM Higgs and the spurion field $\varphi$ are expanded around the VEVs as $H_{1}=(0,v_{H}+h)^{T}/\sqrt{2}$ and $\varphi=v_{\varphi}+\rho$, the dark Higgs $\rho$ can mix with the SM Higgs $h$, so the mass eigenstates become $h_{1}=c_{\alpha}\,h+s_{\alpha}\,\rho$ and $h_{2}=-s_{\alpha}\,h+c_{\alpha}\,\rho$, with $c_{\alpha}\equiv\cos\alpha,s_{\alpha}\equiv\sin\alpha$.
We remark that $S$ and $H_{2}$ do not get VEVs, leaving a $Z_{2}$ symmetry unbroken. $S$, $H_{2}$ and $N_{R,i}$ are odd under the remaining $Z_{2}$, which ensures the stability of the lightest neutral $Z_{2}$-odd particle among them as a dark matter candidate. For small neutrino masses without introducing small parameters, we consider the case where the singlet scalar $S$ is decoupled, so either the lighter neutral component of $H_{2}$ or the lightest RH neutrino can be a dark matter candidate.
III Decoupled scalar and inert doublet scalar masses
From the effective interactions in eq. (2), we can take the limit of integrating out the singlet scalar $S$ whose squared mass becomes $m^{2}_{S}\gg\kappa\kappa^{\prime}$. As a result, we obtain a renormalizable but $Z_{4}$ breaking effective interaction for two Higgs doublets from $\kappa,\kappa^{\prime}$ terms, at tree level as shown in Fig. 1, in the following form,
${\cal L}_{\rm eff}=-\frac{1}{2}\lambda_{5,{\rm eff}}(H^{\dagger}_{1}H_{2})^{2}+{\rm h.c.}$ (3)
with
$\lambda_{5,{\rm eff}}=-\frac{4\kappa\kappa^{\prime}}{m^{2}_{S}}.$ (4)
Here, $\kappa$ is a $Z_{4}$ invariant dimensionful coupling, so it can be of order $m_{S}$. But, $\kappa^{\prime}$ is a $Z_{4}$ breaking coupling due to the VEV of the spurion field $\varphi$, so it is naturally suppressed as compared to $m_{S}$, that is, $\kappa^{\prime}=\lambda^{\prime}_{S\varphi}v_{\varphi}\ll m_{S}$. Therefore, the effective $\lambda_{5,{\rm eff}}$ term in eq. (4) is also suppressed as a consequence of the decoupled scalar.
Figure 1: Feynman diagrams contributing to $\lambda_{5,{\rm eff}}$.
For a small $\lambda_{5}$ term in the effective theory below the threshold of the singlet scalar $S$, we obtain the masses for the components of the inert doublet, $H_{2}=(H^{+},\frac{1}{\sqrt{2}}(H_{0}+iA_{0}))^{T}$, as follows,
$m^{2}_{H^{\pm}}$ $=m^{2}_{2}+\frac{1}{2}\lambda_{3}v^{2}_{H},$ (5)
$m^{2}_{H_{0}}$ $=m^{2}_{2}+\frac{1}{2}\left(\lambda_{L}+\lambda_{5,{\rm eff}}\right)v^{2}_{H},$ (6)
$m^{2}_{A_{0}}$ $=m^{2}_{2}+\frac{1}{2}\left(\lambda_{L}-\lambda_{5,{\rm eff}}\right)v^{2}_{H}$ (7)
where $\lambda_{L}\equiv\lambda_{3}+\lambda_{4}$, with $\lambda_{3},\lambda_{4}$ being mixing quartic couplings in the scalar potential, $\lambda_{3}|H_{1}|^{2}|H_{2}|^{2}+\lambda_{4}(H^{\dagger}_{1}H_{2})(H^{\dagger}_{2}H_{1})$, and $m^{2}_{2}$ is the squared mass parameter for the inert doublet $H_{2}$. Henceforth, we choose $\lambda_{5,{\rm eff}}>0$ without loss of generality such that $A_{0}$ is a dark matter candidate. The most general results for the mass eigenvalues and the mixing angles with the non-decoupled singlet scalar $S$ are shown in eqs. (39), (50) and (51), in the appendix.
In the decoupling limit of the singlet scalar $S$, the small effective $\lambda_{5}$ term gives rise to a mass splitting between two real neutral scalars of $H_{2}$ is suppressed as $\Delta\equiv m_{H_{0}}-m_{A_{0}}\simeq\frac{\lambda_{5,{\rm eff}}v^{2}_{H}}{2m_{0}},$ with $m^{2}_{0}\equiv(m^{2}_{H_{0}}+m^{2}_{A_{0}})/2$. So, we can avoid the strong bound on DM-nucleon scattering with $Z$ boson exchange, as far as $|\Delta|\gtrsim 100\,{\rm keV}$, which sets a lower bound on $\lambda_{5,{\rm eff}}$ as $\lambda_{5,{\rm eff}}\gtrsim 3.3\times 10^{-7}(m_{0}/100\,{\rm GeV})(|\Delta|/100\,{\rm keV})$. Then, for $\kappa\sim m_{S}$, we need $\kappa^{\prime}/m_{S}\gtrsim 8\times 10^{-8}$ for $m_{0}\sim 100\,{\rm GeV}$ and $|\Delta|\gtrsim 100\,{\rm keV}$.
Figure 2: Feynman diagrams contributing to neutrino masses at one-loop.
IV Decoupled scalar and small neutrino masses
In the decoupling limit of the singlet scalar $S$, from the combination of the Yukawa interactions for right-handed neutrinos and a small effective $\lambda_{5,{\rm eff}}$ term in eq. (3), we obtain the one-loop corrections for neutrino masses from the Feynman diagram shown in Fig. 2 as
$({\cal M}_{\nu})_{ij}$ $\simeq$ $\frac{\lambda_{5,{\rm eff}}v^{2}_{H}}{16\pi^{2}}\sum_{k}\frac{y_{N,ik}y_{N,jk}M_{N,k}}{m^{2}_{0}-M^{2}_{N,k}}$ (8)
×$\times\bigg{[}1-\frac{M^{2}_{N,k}}{m^{2}_{0}-M^{2}_{N,k}}\ln\frac{m^{2}_{0}}{M^{2}_{N,k}}\bigg{]}.$
The above results match with those due to a nonzero small $\lambda_{5}$ in the case without a $Z_{4}$ symmetry in Ref. Ma:2006km , and the same results can be obtained from the decoupling limit of the general results in eq. (53) in the appendix. Then, when the RH neutrinos are heavier than the inert doublet Higgs, namely, $M_{N,k}\gtrsim m_{0}$, the above expression with eq. (4) gets simplified to
$({\cal M}_{\nu})_{ij}\simeq\frac{\lambda_{5,{\rm eff}}v^{2}_{H}}{16\pi^{2}}\sum_{k}\frac{y_{N,ik}y_{N,jk}}{M_{N,k}}\bigg{[}\ln\frac{M^{2}_{N,k}}{m^{2}_{0}}-1\bigg{]}.$ (9)
Therefore, for neutrino masses, $m_{\nu}\lesssim 0.1\,{\rm eV}$, choosing $y_{N,ij}={\cal O}(1)$ and $M_{N,k}\sim 10^{9}\,{\rm GeV}$, we can set the upper bound on $\lambda_{5,{\rm eff}}$ as $\lambda_{5,{\rm eff}}\lesssim 8\times 10^{-6}$, which is consistent with the lower bound on $\Delta$ from the direct detection. In other words, we need $\kappa^{\prime}/m_{S}\lesssim 10^{-6}$ for $\kappa\sim m_{S}$ and $M_{N,k}\lesssim\kappa^{\prime}\ll m_{S}$ for $\lambda_{N,kk}\lesssim\lambda^{\prime}_{S\varphi}$. Then, we can keep the RH right-handed neutrinos in the effective theory while the singlet scalar $S$ is decoupled. So, for a small $\lambda_{5,{\rm eff}}$, it is sufficient to take $m_{S}\gtrsim M_{N,k}\sim 10^{9}\,{\rm GeV}$, unless we tune the Yukawa couplings $y_{N,ij}$ to small values.
In the case with $m_{0}\gtrsim M_{N,k}$, the lightest RH neutrino is a dark matter candidate Ibarra:2016dlb whereas the scalars of the inert doublet decay into a pair of the SM lepton and the RH neutrino through the Yukawa interactions. In this limit, eq. (8) becomes approximated to $({\cal M}_{\nu})_{ij}\simeq\frac{\lambda_{5,{\rm eff}}v^{2}_{H}}{16\pi^{2}m^{2}_{0}}\sum_{k}y_{N,ik}y_{N,jk}M_{N,k}.$ For instance, for $m_{\nu}\lesssim 0.1\,{\rm eV}$, $y_{N,ij}={\cal O}(1)$, $m_{0}\sim 100\,{\rm GeV}$ and $M_{N,k}\sim 10\,{\rm MeV}$, we can set the upper bound, $\lambda_{5,{\rm eff}}\lesssim 8\times 10^{-10}$ or $|\Delta|\lesssim 0.2\,{\rm keV}$. But, there is no more lower bound on $\lambda_{5,{\rm eff}}$ because the scalars in the inert doublet decay. A small effective $\lambda_{5,{\rm eff}}$ is achievable if $m_{S}\gtrsim 10^{9}\,{\rm GeV}(\lambda^{\prime}_{S\varphi}/\lambda_{N,kk})$ for $v_{\varphi}=10\,{\rm MeV}/\lambda_{N,kk}$ for $\kappa\sim m_{S}$.
In Fig. 3, we show the parameter space for inert doublet scalar mass $m_{0}$ vs RH neutrino mass $M_{N}\equiv M_{N,1}$ satisfying $m_{\nu}=0.1\,{\rm eV}$. The color code indicates a scan over $\lambda_{5,{\rm eff}}$ between $10^{-8}$ and $10^{-4}$, showing that $\lambda_{5,{\rm eff}}\gtrsim 10^{-6}$ and $M_{N}\gtrsim 2\times 10^{7}\,{\rm GeV}$ are necessary for $m_{0}\gtrsim 100\,{\rm GeV}$. In Fig. 4, we decode the correlation between neutrino mass $m_{\nu}$ and singlet scalar mass $m_{S}$ for varying $M_{N}$ between $10^{4}\,{\rm GeV}$ and $10^{9}\,{\rm GeV}$. In both Figs. 3 and 4, we imposed $|\Delta|>100\,{\rm keV}$ and took into account the perturbativity, $|\lambda_{5,{\rm eff}}|<1$. Then, we need the singlet scalar mass $m_{S}$ to be of order $10^{14}-10^{16}\,{\rm GeV}$ for $m_{\nu}=0.01-0.1\,{\rm eV}$, which might hint at a connection of neutrino mass generation to the unification. If we choose smaller Yukawa couplings $y_{N}$, the diagonal boundary in Fig. 3 is shifted to the left, allowing for smaller $M_{N}$ for a fixed $m_{0}$, and the strip in Fig. 4 is shifted to the bottom, making $m_{S}$ smaller for a fixed $m_{\nu}$.
Figure 3: Inert doublet scalar mass $m_{0}$ vs RH neutrino mass $M_{N}$ for varying $\lambda_{5,{\rm eff}}$.
Figure 4: Neutrino mass $m_{\nu}$ vs singlet scalar mass $m_{S}$ for varying $M_{N}$
V Dark matter direct detection
When the lighter neutral scalar from the inert doublet is dark matter, we can avoid the strong bound from direct detection if the mass splitting between $A_{0}$ and $H_{0}$ is large enough, i.e. $|\Delta|\gtrsim 100\,{\rm keV}$. Then, the cross section for spin-independent scattering between dark matter and nucleus is dominated by the tree-level exchanges of Higgs-like scalars Barbieri:2006dq ; LopezHonorez:2006gr , as follows,
$\sigma^{\rm SI}_{A_{0}}=\frac{\lambda_{L5}^{2}\mu^{2}_{N}(Zf_{p}+(A-Z)f_{n})^{2}}{4\pi m^{2}_{A_{0}}A^{2}}\bigg{(}\frac{c^{2}_{\alpha}}{m^{2}_{h_{1}}}-\frac{s^{2}_{\alpha}}{m^{2}_{h_{2}}}\bigg{)}^{2}$ (10)
where $\lambda_{L5}\equiv\lambda_{L}-\lambda_{5,\,{\rm eff}}$, $m_{h_{1,2}}$ are the masses for SM Higgs-like and dark Higgs-like scalars, $Z,A-Z$ are the numbers of protons and neutrons in the detector nucleus, $\mu_{N}$ is the reduced mass for the DM-nucleus system, and $f_{p,n}$ are the nucleon form factors Hoferichter:2015dsa ; Junnarkar:2013ac ; Kim:2023pwf .
The DM relic density is determined by the DM annihilations into a pair of SM fermions, $WW,ZZ$ and $h_{i}h_{j}$ with $i,j=1,2$ Barbieri:2006dq ; LopezHonorez:2006gr ; Gong:2012ri . If $H_{0},H^{\pm}$ have comparable masses to $m_{A_{0}}$, we also need to include the co-annihilation channels associated with them as eq. (57) in the appendix. In particular, there is room to adjust $\lambda_{L}$ and dark Higgs-portal couplings, denoted by ${\cal L}_{\rm int}\supset-\frac{1}{2}\sum_{i=1,2}\lambda_{H_{i}\varphi}|H_{i}|^{2}\varphi^{2}$, to explain the DM relic density for relatively heavy DM due to $A_{0}A_{0}\to h_{i}h_{j}$ while the direct detection bound is satisfied Belyaev:2016lok . DM masses above about $100\,{\rm GeV}$ are consistent with the bounds from direct detection, as far as $|\lambda_{L5}(c^{2}_{\alpha}-s^{2}_{\alpha}m^{2}_{h_{1}}/m^{2}_{h_{2}})|\lesssim 10^{-2}$, apart from the resonance, and the bounds from the SM Higgs decays can be easily satisfied for $m_{A_{0}},m_{h_{2}}\gtrsim m_{h_{1}}/2=62.5\,{\rm GeV}$. But, for a smaller value of $|\lambda_{L5}|$, as required near resonances, the one-loop corrections to the DM-nucleus cross section are relevant Abe:2015rja .
In Fig. 5, we depict the DM relic density as a function of DM mass for $\lambda_{H_{2}\varphi}=0,0.01,0.03$, in red solid, blue dashed and dotted lines, respectively. We chose a zero Higgs mixing, $\alpha=0$, the singlet parameters are $\lambda_{H_{1}\varphi}=0$, $v_{\varphi}=10^{4}\,{\rm GeV}$ and $m_{h_{2}}=m_{\rho}=80\,{\rm GeV}$, and the parameters for the inert doublet are $\lambda_{L5}=0.01$, $\Delta=10\,{\rm MeV}$ and $m_{H^{\pm}}=m_{0}+1\,{\rm GeV}$. Thus, as $\lambda_{H_{2}\varphi}$ increases, the correct relic density can be explained for a heavier DM mass far beyond $100\,{\rm GeV}$. The results are insensitive to $\Delta$ as far as $\Delta\lesssim T_{f}$ where $T_{f}$ is the freeze-out temperature given by $T_{f}\simeq m_{0}/20$. Moreover, we remark that a nonzero $\lambda_{H_{1}\varphi}$ also contributes to the dark matter annihilation into $\varphi\varphi$ with the SM Higgs exchange, but its contribution is suppressed for a small $|\lambda_{L5}|$.
Figure 5: DM relic density as a function of DM mass $m_{0}$, depending on the inert doublet portal coupling, $\lambda_{H_{2}\varphi}$.
VI Thermal leptogenesis
Baryogenesis can be realized in our model via thermal leptogenesis Fukugita:1986hr , due to the decays of RH neutrinos with hierarchical masses. The RH neutrinos can be once in thermal equilibrium with the SM plasma in the early universe due to their Yukawa couplings to the inert doublet or the spurion scalar, which mixes with the SM Higgs. As the asymmetries generated from the decays of the heavier RH neutrinos, $N_{2,3}$, are negligible due to strong washout effects by scattering, we focus on the $N_{1}$ contribution to the lepton asymmetry. The baryon-to-photon ratio converted from the final $B-L$ asymmetry is given by
$\eta_{B}=-C\varepsilon_{1}\kappa_{1}$ (11)
where $C\simeq 0.01$, $\varepsilon_{1}$ is the $CP$ asymmetry parameter, and $\kappa_{1}$ is the efficiency factor approximated Buchmuller:2004nz to
$\kappa_{1}(K_{1})\simeq\frac{2}{z_{B}(K_{1})K_{1}}\Big{(}1-e^{-\frac{1}{2}z_{B}(K_{1})K_{1}}\Big{)},$ (12)
with $K_{1}=\Gamma_{1}/H(z_{1}=1)$ being the decay parameter and $z_{B}$ being the value of $z_{1}=M_{N,1}/T$ at the completion of the baryon asymmetry generation, given Buchmuller:2004nz by
$z_{B}(K)\simeq 1+\frac{1}{2}\ln\bigg{(}1+\frac{\pi K^{2}}{1024}\bigg{[}\ln\Big{(}\frac{3125\pi K^{2}}{1024}\Big{)}\bigg{]}^{5}\bigg{)}.$ (13)
Here, $\Gamma_{1}$ is the decay rate for $N_{1}$, given by $\Gamma_{1}\simeq\frac{M_{N,1}}{8\pi}(y^{\dagger}_{N}y_{N})_{11}$ for $m_{0}\ll M_{N,1}$, and $H(z_{1})$ is the Hubble parameter. In the Casas-Ibarra parametrization for the Yukawa matrix $y_{N}$ Casas:2001sr , we can express $(y^{\dagger}_{N}y_{N})_{ij}=\sqrt{\Lambda_{i}\Lambda_{j}}{\widetilde{m}}_{ij}$ in terms of the effective neutrino mass matrix ${\widetilde{m}}\equiv RD_{{\cal M}_{\nu}}R^{\dagger}$ with $R$ being arbitrary complex rotation matrix, $D_{{\cal M}_{\nu}}={\rm diag}(m_{\nu_{1}},m_{\nu_{2}},m_{\nu_{3}})$, and $\Lambda_{i}=\frac{4\pi^{2}\xi_{i}M_{N,i}}{\lambda_{5,{\rm eff}}v^{2}_{H}}$ Hugle:2018qbw . Here, $\xi_{i}=\big{(}\frac{1}{8}\frac{M^{2}_{N,i}}{m^{2}_{H_{0}}-m^{2}_{A_{0}}}[L(m^{2}_{H_{0}})-L(m^{2}_{A_{0}})]\big{)}^{-1}$ with $L(m^{2})=\frac{m^{2}}{m^{2}-M^{2}_{N,i}}\ln(m^{2}/M^{2}_{N,i})$ are order one parameters in most of the parameter space. Thus, we get the decay parameter as $K_{1}\simeq\frac{2\pi^{2}\xi_{1}}{\lambda_{5\,{\rm eff}}}\sqrt{\frac{45}{64\pi^{5}g_{*}}}\frac{M_{P}}{v^{2}_{H}}{\widetilde{m}}_{11}$ for $m_{0}\ll M_{N,1}$.
Figure 6: Parameter space for RH neutrino mass $M_{N}$ vs $\lambda_{5,{\rm eff}}$, showing $\eta_{B}>6.1\times 10^{-10}$ in purple region. The contour lines show the parameter space for the heaviest neutrino mass equal to $m_{\nu}=0.05\,{\rm eV}$.
For the case with three RH neutrinos, the $CP$ asymmetry parameter is bounded Davidson:2002qv ; Hugle:2018qbw by
$\varepsilon_{1}\lesssim\frac{3\pi}{4\lambda_{5\,{\rm eff}}v^{2}_{H}}\,\xi_{3}(m_{h}-m_{l})M_{N,1}$ (14)
with $m_{h},m_{l}$ are the masses for the heaviest and lightest active neutrinos, respectively. After optimizing the efficiency factor $\kappa_{1}$ and the $CP$ asymmetry parameter $\varepsilon_{1}$, we obtain the relation, $\lambda_{5,{\rm eff}}\simeq 4\pi\xi_{1}(m_{l}/10^{-3}\,{\rm eV})$ Hugle:2018qbw , for which the baryon-to-photon ratio without washout effects from $\Delta L$ processes is given by
$\eta_{B}\simeq 3.8\times 10^{-10}\Big{(}\frac{M_{N,1}}{10^{4}\,{\rm GeV}}\Big{)}\bigg{(}\frac{10^{-4}}{\lambda_{5,{\rm eff}}}\bigg{)}\bigg{(}\frac{m_{h}}{0.1\,{\rm eV}}\bigg{)}.$ (15)
[…]
We remark that the asymmetry in the inert doublet or the DM asymmetry can reduce the lepton asymmetry, because $lH_{2}\to{\bar{l}}{\bar{H}}_{2}$ is favored over ${\bar{l}}{\bar{H}_{2}}\to lH_{2}$ Racker:2014yfa . However, for a sizable $\lambda_{5,{\rm eff}}$, $H_{2}{\bar{H}}_{1}\leftrightarrow{\bar{H}}_{2}H_{1}$ process is fast, so it can make the DM asymmetry Boltzmann sufficiently suppressed. It turns out that $\Delta L=2$ processes are inefficient if $|\lambda_{5,{\rm eff}}|\gtrsim 10^{-5}$ for $M_{N,1}\lesssim 10^{9}\,{\rm GeV}$ Clarke:2015hta .
VII Conclusions
We presented a unified solution with $Z_{4}$ discrete symmetry for small neutrino masses and stability of dark matter in the radiative seesaw scenarios. Once the $Z_{4}$ symmetry is broken to $Z_{2}$ by the VEV of the spurion singlet scalar $\varphi$ much below the mass of the singlet scalar $S$, a small effective lepton number violating interaction between two Higgs doublets is generated, being responsible for a small mass splitting of the neutral inert scalars at tree level and small neutrino masses at one-loop level.
Although the $Z_{2}$ symmetry is sufficient for stability of dark matter, its extension to the $Z_{4}$ symmetry in our model allows for explaining the origin of small neutrino masses, even if inert doublet scalar and RH neutrinos are relatively light and the Yukawa couplings for neutrinos are sizable. This option gives rise to phenomenologically a novel window for new particle searches. Moreover, we showed that the spurion scalar responsible for the $Z_{4}$ symmetry breaking opens up a new parameter space for obtaining the correct relic density for inert doublet-like dark matter and a successful leptogenesis is achieved.
Acknowledgments
The work is supported in part by Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education, Science and Technology (NRF-2022R1A2C2003567 and RS-2024-00341419 (JKK)). This research was supported by the Chung-Ang University research grant in 2019.